zackariahharriett:

An unknown number y is 15 more than an unknown number x. The number y is also x less than 8. The equations to find x and y are shown below. y = x + 15 y = −x + 8 Which of the following statements is a correct step to find x and y? (4 points) Multiply the equations to eliminate y. Add the equations to eliminate x. Write the points where the graphs of the equations intersect the x-axis. Write the points where the graphs of the equations intersect the y-axis.

b3lla2025:

what ones can you eliminate

AZ:

I asked you in my reply to your earlier question but did not receive a response, so I'm going to assume that you don't know the substitution or elimination method. Substitution is when you take one equation and substitute it into the other equation to solve for one variable. So for example, we have 2a + 3b = 30 a + 2b = 15 we can manipulate the second equation a + 2b = 15 and get a = -2b + 15 we can then take this equation and SUBSTITUTE it into the first equation in place of the 'a' there 2a + 3b = 30 2(-2b+15) + 3b = 30 And then we can solve for b. Once we find b, we can then plug it into either equation and solve for 'a'. Now that's the substitution method.

AZ:

The elimination method is when you manipulate both equations so that way you're able to ELIMINATE one of the variables so that you can solve for remaining variable. Here's an example: a + b = 40 3a - 2b = 20 If we multiply the first equation by 2 on both sides, we get 2(a+b) = 2(40) 2a + 2b = 80 now we have 2a + 2b = 80 3a - 2b = 20 If we add both the equations we get (2a+2b) + (3a-2b) = 80 + 20 2a + 2b + 3a - 2b = 80 + 20 5a + 2b - 2b = 80 + 20 5a = 100 Do you see how we eliminated the variable 'b' and were able to solve for a? After you find 'a', you can then plug it into either equation and solve for b. That's the elimination method.

AZ:

To answer your question, what do we have to do to solve for x and y? The solution (x, y) is going to be where the two lines intersect each other. Knowing where the equations cross either the x-axis or y-axis doesn't tell us what the solution is. It might help us in graphing the equations to find where they intersect but it's not the correct step to find x and y.

AZ:

So if we want to use the elimination method to eliminate y, what would we do? If we want to eliminate x, what do we do? They don't have any number (or coefficient) in front of it so that means you don't have to multiply both sides by a number to get the same number so you can eliminate it.

zackariahharriett:

oh my bad guys i was away from my computer

zackariahharriett:

i did awswer your question i was doing it in the process which made me get my previous question right

AZ:

so did you get the answer for this question as well?